OpenStudy (anonymous):

Given the linear equation y= 1/3x-3 find the y-coordinates of the points (-6,_)(-3,_)and (6,_)

OpenStudy (angela210793):

just put the numbers u're given and u'll find the one which is missing, for example y=x/3-3 x=-6 y=-6/3-3=-3-3=-6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(x,y) is a coordinate point

OpenStudy (anonymous):

you are given the x find the y

OpenStudy (anonymous):

i know but how

OpenStudy (anonymous):

\[y=\frac{1}{3}x-3\] (-6,_) x = -6 \[y=\frac{1}{3}(-6)-3\]

OpenStudy (anonymous):

know how to multiply fractions?

OpenStudy (anonymous):

its been a few years an im not that good at algebra

OpenStudy (anonymous):

k, a whole number can be written as a fraction over 1 \[\frac{1}{3}(-6)=\frac{1}{3}*\frac{-6}{1}=\frac{-6}{3}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Does that help any?

OpenStudy (anonymous):

yes but do i need to draw up a grapgh i dont know how to find the y coordinate just multiply fractions for ex.-6/1

OpenStudy (anonymous):

The y coordinate is given by the equation \[y=\frac{1}{3}x-3\]

OpenStudy (anonymous):

replace the x with the value given in the coordinates (-6,_) \[y=\frac{1}{3}(-6)-3\] solve that and you have the y value of that coordinate

OpenStudy (anonymous):

how

OpenStudy (anonymous):

plz help me someone

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

! yes

OpenStudy (anonymous):

(-6,-5) is your first coordinate

OpenStudy (anonymous):

thank u but can you set the other 2 up for me then ill wk them

OpenStudy (anonymous):

(-3, ) \[y=\frac{1}{3}(-3)-3\] (6, ) \[y=\frac{1}{3}(6)-3\]

OpenStudy (anonymous):

ok (-3,-4)

OpenStudy (anonymous):

and (6,-1)

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

thank you so much,,im super grateful

OpenStudy (anonymous):

glad to try and help =)

OpenStudy (anonymous):

udid :)

OpenStudy (anonymous):

good work guys